OpenStudy (jbeanz97):

Your friend wants to borrow $1,137.45 from you to pay off a credit card that charges a 14.7% APR. You agree to the loan but require your friend to pay you interest of 3.6% APR on the loan and your friend agrees. Your friend pays you $200.00 at the end of the first month. How much goes toward the principal?

OpenStudy (jbeanz97):

@ganeshie8 @Hero @phi help?

OpenStudy (justin9529):

1) The outstanding balance accumulates with interest like this: $1,137.45 (0.036/12) = $3.41. Now, out of the $200, 200-3.41 = $196.59 goes toward the principal Her friend would have paid interest = $1,137.45 (0.147/12) = $13.93. So she saved $13.93 - $3.41 = $10.52 Hope that helps

OpenStudy (jbeanz97):

thanks! could you help with the second part to that question?

OpenStudy (justin9529):

yea sure

OpenStudy (jbeanz97):

2. how much interest does your friend save compared to the credit card at the end of the first month?

OpenStudy (justin9529):

shoot im actually im not sure:/

OpenStudy (jbeanz97):

oh wait would the answer to number 2 be $10.52?

OpenStudy (wolf1728):

Justin - you and jbean in the same class?

OpenStudy (justin9529):

im a junior

OpenStudy (jbeanz97):

im a junior too but im taking consumers math this year

OpenStudy (wolf1728):

By that I mean you answered "part 2 " of a question that wasn't even asked.

OpenStudy (justin9529):

lol i looked it up real quick

OpenStudy (wolf1728):

looked it up quick where? You have access to questions not everyone can even see?

OpenStudy (justin9529):

take your question and copy and past into google

OpenStudy (justin9529):

paste*

OpenStudy (wolf1728):

well at least it is nice to know how people are getting the questions and answers

OpenStudy (wolf1728):

For this problem wouldn't it be nice to know how many years these payments will be made? (Shouldn't that be the first step?)

OpenStudy (wolf1728):

I calculate .479 years. Agreed?

OpenStudy (wolf1728):

Do you folks have any formulas for this? These formulas are pretty difficult: Here is the formula for years (months actually) http://www.1728.org/loanfrm3.htm Guess you folks left.

OpenStudy (justin9529):

holy poop no lol